### v1

@@ Abstract
Text-to-Image (T2I) Diffusion Models (DMs) have shown impressive abilities in generating high-quality images based on simple text descriptions. However, as is common with many Deep Learning (DL) models, DMs are subject to a lack of robustness. While there are attempts to evaluate the robustness of T2I DMs as a binary or worst-case problem, they cannot answer how robust in general the model is whenever an adversarial example (AE) can be found. In this study, we first introduce a probabilistic notion of T2I DMs’ robustness; and then establish an efficient framework, ProTIP, to evaluate it with statistical guarantees. The main challenges stem from: i) the high computational cost of the generation process; and ii) determining if a perturbed input is an AE involves comparing two output distributions, which is fundamentally harder compared to other DL tasks like classification where an AE is identified upon misprediction of labels. To tackle the challenges, we employ sequential analysis with efficacy and futility early stopping rules in the statistical testing for identifying AEs, and adaptive concentration inequalities to dynamically determine the “just-right” number of stochastic perturbations whenever the verification target is met. Empirical experiments validate the effectiveness and efficiency of ProTIP over three common T2I DMs. Finally, we demonstrate an application of ProTIP to rank commonly used defence methods.
@@ 1 Introduction
Text-to-Image (T2I) Diffusion Models (DMs) have made impressive strides in recent years. State-of-the-art (SOTA) models, like DALL-E 3[3], Imagen [43], and Parti [58] can generate high-quality images from textual descriptions. Additionally, Stable Diffusion [41] represents a successful large-scale implementation of the latent diffusion model. With the advent of T2I DMs, we can now create visually stunning images with simple text prompts.
@@
However, studies [65, 10, 11] have demonstrated that small perturbations to the textual inputs can substantially degrade the performance of T2I DMs. Fig. 1 illustrates examples where introducing minor perturbations to the prompt leads to a change in the generated image outcome, which raises concerns about the model’s robustness in the context of its downstream applications [5, 30]. A pivotal question arises: how can we systematically evaluate and verify (when a specific verification target is provided) the robustness of T2I DMs?
@@
Figure 1: Examples illustrating perturbations applied to the prompt for Stable Diffusion, employing two methods as described in Sec. 3.2
@@
The problem of lacking robustness in T2I DMs is not unexpected, considering that robustness issues are pervasive in the broader realm of Deep Learning (DL) [45, 14]. Generally, robustness is defined as the invariant decision of the DL model against small perturbations on inputs. A small perturbation on an input is termed as an Adversarial Example (AE) if it leads to a different prediction from the label assigned to the original input. Over the past decade, numerous studies have attempted to frame DL robustness verification as a binary or worst-case problem, addressing questions like “if AEs exist (in a small local region of the input)?” or “what is the closest AE to the input?” [59, 6, 21]. Emerging studies are adopting a probabilistic view, formulating robustness verification as a statistical inference problem [53, 48, 62, 47, 20, 8, 7]. Such probabilistic robustness notion is arguably of more practical interest than binary/worst-case ones, because it provides an overall evaluation of how robust the model is whenever an AE can be found [53, 20] and accepts residual risks that are more realistic to achieve [48, 62]. We concur with this view and argue that T2I DMs also necessitate probabilistic robustness verification, which, to the best of our knowledge, is absent in SOTA.
@@
In this work, we first formally define the probabilistic robustness of T2I DMs against character-level stochastic perturbation; then, we devise an efficient framework, named ProTIP, to evaluate and verify it (assuming a verification target is given). In addressing the statistical inference problems within ProTIP, we encounter new challenges specific to the generative nature of DMs. These challenges, distinct from traditional DL tasks like classification or regression, include: i) the generation process is very computationally insensitive; ii) the identification of a perturbed input as an AE entails comparing distributional differences, which is fundamentally harder compared to DL tasks like classification where an AE is identified by a change of the prediction label. To tackle the challenges, we employ sequential analysis with efficacy and futility early stopping rules in the statistical testing for indicating AEs, and adaptive concentration inequalities to dynamically determine the “just-right” number of stochastic perturbations whenever the verification target is met. Experimental results validate the effectiveness and efficiency of our ProTIP, while an application of ProTIP is also presented to showcase its use in ranking common defences against text perturbations. In summary, key contributions of this paper include:
@@
  • Problem formulation: For the first time, we formulate the probabilistic robustness verification problem for T2I DMs against stochastic perturbation, with formalised evaluation metrics.
@@
  • Efficient solution: To solve the formulated problem of probabilistic robustness verification for T2I DMs, we have developed an efficient framework, ProTIP, which incorporates several sequential analysis methods to dynamically determine the sample size and thus enhance ProTIP’s efficiency.
@@ 2.1 Text-to-Image Diffusion Models
Generative AI has been thriving in the multi-modal field, with diffusion models emerging as a powerful new family of deep generative models with record-breaking performance in many applications, including video generation [56], image reconstruction [46], and T2I generation [3]. DMs are a class of probabilistic generative models that apply a noise injection process, followed by a reverse procedure for sample generation. Importantly, DMs allow a notion of control at generation time through so-called guidance, enabling several complicated generative tasks, among which T2I DMs are examples of guided image generation tasks using descriptive text as the guidance signal. Models like OpenAI’s Stable Diffusion [41] and Google’s Imagen [43], trained on large-scale datasets of annotated text-image pairs, can generate high-quality images based on simple natural language descriptions. In this context, Glide [35] emerges as a text-guided diffusion model that not only supports image generation but also excels in the realm of image editing. Moreover, commercial products such as DALL-E 2 and Midjourney [1] have showcased stunning effects across various scenarios, contributing to the richness and diversity of text-to-image applications. We refer readers to Appendix 0.A for more details on T2I DMs.
@@ 2.2 Deep Learning Robustness
DL models, such as Neural Networks, are notoriously unrobust to small perturbations [59, 6, 21]. While definitions of robustness vary in literature, they share a common intuition that a DL model’s decision should remain invariant against small perturbations on a given input—typically it is defined as all inputs in a region $\eta$ have the same prediction label, where $\eta$ is a small norm ball (in a $L_{p}$-norm distance) of radius $\gamma$ around an input $x$ (normally called the test seed). A perturbed input (e.g., by adding noise on $x$) $x^{\prime}$ within $\eta$ is an AE if its prediction label differs from $x$.
In the Safe AI community, DL robustness has been the property in the spotlight. Many studies on evaluating the DL robustness have been done, and they frame the problem in different ways. In Fig. 2, we summarise 4 common ways of formulating the problem [8]. Earlier works, such as [13, 23, 42], formulate the verification problem as a binary question by asking if any AEs can be found within a given input norm-ball of a specified radius, cf. Fig. 2(a). Such “binary robustness” can be normally evaluated in two ways [23, 42, 57]: either through reachability algorithms that aim to determine the lower and upper bounds of the output within the input norm-ball $\eta$, using layer-by-layer analysis; or by solving it using SAT/SMT solvers as a variety of constraint-based programming problems. Fig. 2(b) poses a similar yet different question: what is the maximum radius of $\eta$ such that no AE exists within it? Intuitively, it is finding the “largest safe distance” for the test seed $x$ [2, 33, 54, 55]. In Fig. 2(c), on the other hand, it evaluates the model’s robustness by introducing adversarial attacks to cause the maximum prediction loss within the specified norm ball $\eta$. This method is often applied in adversarial training to enhance the model’s robustness to resist attacks [31, 50].
@@
The three aforementioned methods all estimate the robustness of the model by detecting the presence of AE or the AE that gives the maximum loss/safe-radius. This type can be called deterministic robustness [62]. As argued by [53], the three aforementioned methods suffer from two major drawbacks: i) they fail to convey how robust the model is whenever an AE is found; ii) they pose scalability challenges when the model is large. To overcome these limitations, more recent works [53, 48, 62, 47, 20, 8, 7] develop a new probabilistic view, by defining robustness as the proportion of AEs inside the norm-ball $\eta$, cf. Fig. 2(d). This probabilistic metric is arguably more practical for two reasons: i) binary/worst-case robustness focusing on extreme cases is neither necessary nor realistic, especially when the model is large; what is more relevant is knowing the proportion of AEs; ii) since all practical applications have an acceptable level of risk, it suffices to demonstrate that the probability of a violation is below a required threshold, rather than confirming it to be exactly zero. Without loss of generality, we illustrate probabilistic robustness of a DL classification task as:
@@ 2.3 Robustness of Text-to-Image Diffusion Models
Despite variations of definitions, the key to evaluating and improving the model’s robustness is detecting AEs. The approaches to this end are often referred to as “adversarial attacks” (i.e., via input perturbations). For T2I models, adversarial attack methods can also be classified into character-level, word-level, sentence-level or multi-level, depending on the granularity of input perturbations.
Recent studies [32, 10] have shown that T2I DMs are very sensitive to black-box attacks like text perturbation. Although such perturbations are commonly referred to as “attacks”, they are not necessarily malicious actions from attackers [64]. They may also represent natural human errors that follow a stochastic generation process. In [11], it proved that T2I DMs are vulnerable to realistic errors (e.g., typo, glyph, phonetic) that humans can make and reveal their non-trivial robustness issues. The work in [65] revealed the vulnerability of T2I models and attributes it to the weak robustness of the used text encoders. Aforementioned existing works only revealed the lack of robustness of T2I DMs in the face of text perturbations and focused on deterministic robustness (e.g., Maximised Prediction Loss). To the best of our knowledge, there exists no dedicated exploration into the probabilistic robustness of T2I models when they subject to stochastic perturbations, and our ProTIP is the first verification framework for this problem, backed by statistical guarantees. Moreover, ProTIP addresses unique challenges arising from the generative characteristics of T2I DMs, by adopting sequential analysis and adaptive concentration inequality to improve the verification efficiency, details of which are provided in the following section.
@@ 3.1 Problem Statement
where $T$ denotes the CLIP [40] model’s text encoder, $s$ is a similarity measurement function (e.g., cosine similarity), $\gamma$ denotes a given threshold on similarity. While $I$ is an indication function as defined in Def. 1, its value now depends on whether the output distributions before and after the perturbation differ. $Pr(x^{\prime})$ indicates the probability that $x^{\prime}$ is the next perturbed text generated.
@@ 3.2 Stochastic Perturbation Generation
While ProTIP can cope with any text perturbations generated stochastically, without loss of generality (WLOG), we study five character-level text perturbation methods from [34]: insert, substitute, swap, delete, and keyboard error. First, we set the perturbation rate like 10%, indicating that 10% of the words in the sentence will be perturbed. Then, one of the six perturbation methods as shown in Table 1 will be randomly selected. Taking insertion perturbation as an example, a random letter is chosen from the 26 English alphabets. Subsequently, a random word in the sentence is selected, and a random position within that word is chosen to insert the selected letter. Next, check if the 10% perturbation rate is met by the perturbed sentence. If not, continue the perturbation process until we meet the perturbation rate requirement; if yes and not a duplicate, stop perturbation. Table 1 lists the details and examples of perturbations operations given the sample input “A red ball on green grass under a blue sky”.
Table 1: Examples of stochastic text perturbations
Perturbation
Description
Example
Insert
Insert a character randomly
A reed ball on green grass under a blue sky.
Substitute
Substitute a character randomly
A red ball on green grass under a blae sky.
Swap
Swap two characters randomly
A red ball on green garss under a blue sky.
Delete
Delete a character randomly
A red ball on reen grass under a blue sky.
Keyboard
Substitute a char. by keyboard distance
A red ball on green grass under blue sjy.
@@
As Def. 1, the intuition behind robustness suggests that the outputs of a DL model should remain invariant to perturbations on the inputs, while the inputs, before and after the perturbation, should be deemed very similar—thus, Def. 1 restricts the perturbation in a norm-ball $B(x,\gamma)$ with a small radius $\gamma$. Such perturbation distance control is easily achievable in Computer Vision where the inputs (either raw pixel values or the latent space feature values) can be normalised and projected into a continuous input space. However, for DL models take text inputs (like T2I DMs), the discrete nature of text make it more challenging to study robustness [27] as a single character change may completely alter the semantics meaning. E.g., “a red ball on …” and “a red bell on …” have totally different semantic meanings with only one character perturbed. Thus, to study the robustness of T2I DMs, first we need to ensure that the perturbed text $x^{\prime}$ has the same semantics to the original text $x$. Following the idea from [11, 65, 29, 60] for the same problem, we employ CLIP [40] as the model for text encoding, leveraging CLIP’s ability to represent image and text information while preserving their relationships. That is, as per Def. 2, we use CLIP’s text encoder $T$ to extract the original embeddings $T(x)$ and perturbed sentence embeddings $T(x^{\prime})$ to compute the similarity score $s(T(x),T(x^{\prime}))$ (e.g., by cosine similarity). Only when the similarity score is greater than a given threshold $\gamma$, we deem $x^{\prime}$ is a valid perturbation, where $\gamma$ can be determined by the method developed in [27] (which yields the maximum perturbation distance in the continuous embedding space that preserves the semantics meaning for a given text). In this way, we ensure that only valid perturbations (sharing similar semantics as the original text) are used for robustness evaluation in ProTIP.
@@ 3.3 Indication of Adversarial Examples
which corresponds to the cosine similarity between visual CLIP embedding $V(y)$ for an image $y$ and textual CLIP embedding $T(x)$ for a caption $x$. The score is bound between 0 and 100, and the closer to 100, the better. Then, for each image $y^{\prime}_{i}$, we can calculate the “CLIP score” $c^{\prime}_{i}$ as the coexistence likelihood of the image $y^{\prime}_{i}$ and the original text $x$, by measuring $c^{\prime}_{i}:=CLIPScore(x,y_{i}^{\prime})$, yielding a set of i.i.d. CLIP scores $\{c^{\prime}_{1},\dots,c^{\prime}_{n}\}$. We can repeat the same calculation for each generated image $y_{i}$ from the original text $x$, and collectively form another set of i.i.d. CLIP scores $\{c_{1},\dots,c_{n}\}$ corresponding to $x$.
@@
  • $\mathcal{H}_{1}$: The CLIP scores of the $c^{\prime}_{i}$ group is smaller than the $c_{i}$ group
While there are established statistical testing methods, e.g., the t-test, u-test, and their variants [25], that can be employed to answer the above question, their applicability depends on whether the data distribution meets certain assumptions. Our framework ProTIP is compatible with any statistical tests in this step, cf. later Sec. 4 for our choice in the experiments. We can now implement the AE indicator function in Def. 2 as:
@@
In contrast to conventional DL tasks like classification, where detecting an AE involves an “one-off” test to see if the prediction label changes, determining if a perturbation affects the T2I DM requires multiple generations and comparing the distributional differences between two sets of images generated before and after the perturbation. This shift in the evaluation approach leads to a significantly increased computational workload. To improve the efficiency of solving the hypothetical testing in Eq. (5), we employ sequential analysis with early stopping rules [22, 38, 51]. Instead of collecting all data with the maximum sample size (fixed at the design time) and then analysing the data a single time by the end, sequential analysis conducts interim analyses during data collection, so that we can prematurely stop data collection at an interim analysis upon rejecting or accepting the null hypothesis. Thanks to early stopping, sequential designs will, on average, require fewer samples [25, Chap. 10].
In ProTIP, we adopt sequential analysis with both Efficacy and Futility stopping rules, by controlling the $\alpha$ (Type I error, false positive when a null hypothesis is incorrectly rejected) and the $\beta$ (Type II error, false negative when the null hypothesis is accepted and it is actually false), respectively. Intuitively, the Efficacy stopping rule says, if the analysis reveals a statistically significant result at some interim stage, data collection can be terminated because we reject $\mathcal{H}_{0}$ when observing the p-value $p<\alpha_{k}$ (where $\alpha_{k}$ is the cumulative Type I error rate at the $k$-th interim analysis, controlled by alpha-spending functions [25]). On the other hand, we may also stop for futility, which means it is either impossible or very unlikely for the final analysis to yield $p<\alpha$. This can be implemented by controlling the Type II error rate across interim analysis using a beta-spending function. We refer readers to [25, Chap. 10] for more details on the topic.
Specifically, we employ the R package rpact [52] to perform the sequential analysis by setting Pocock type alpha/beta spending functions [15] to determine the sample size and the thresholds for efficacy/futility early stopping rules. Details are presented in Sec. 4 and Appendix 0.C.

### v2

@@ Abstract
Text-to-Image (T2I) Diffusion Models (DMs) excel at creating high-quality images from text descriptions but, like many deep learning models, suffer from robustness issues. While there are attempts to evaluate the robustness of T2I DMs as a binary or worst-case problem, they cannot answer how robust in general the model is whenever an adversarial example (AE) can be found. In this study, we first formalise a probabilistic notion of T2I DMs’ robustness; and then devise an efficient framework, ProTIP, to evaluate it with statistical guarantees. The main challenges stem from: i) the high computational cost of the image generation process; and ii) identifying if a perturbed input is an AE involves comparing two output distributions, which is fundamentally harder compared to other DL tasks like classification where an AE is identified upon misprediction of labels. To tackle the challenges, we employ sequential analysis with efficacy and futility early stopping rules in the statistical testing for identifying AEs, and adaptive concentration inequalities to dynamically determine the “just-right” number of stochastic perturbations whenever the verification target is met. Empirical experiments validate ProTIP’s effectiveness and efficiency, and showcase its application in ranking common defence methods.
@@ 1 Introduction
Recent advancements in Text-to-Image (T2I) Diffusion Models (DMs), including state-of-the-arts (SOTA) like DALL-E 3 [3], Imagen [44], Parti [61] and Stable Diffusion [42], enable the generation of high-quality images from text prompts. However, studies [69, 10, 11] have shown that small perturbations in textual input can substantially degrade the performance of T2I DMs. Fig. 1 illustrates examples in which introducing a minor perturbation to the prompt leads to a non-trivial change in the generated images, raising concerns about the model’s robustness in the context of its downstream applications [5, 32]. As such, a pivotal question arises: how can we systematically evaluate and verify (when a specific verification target is provided) the robustness of T2I DMs?
@@
Figure 1: Examples illustrating perturbations applied to the prompt for Stable Diffusion.
@@
The problem of lacking robustness in T2I DMs is not unexpected, considering that robustness issues are pervasive in the broader realm of Deep Learning (DL) [48, 14]. Generally, robustness is defined as the invariant decision of the DL model against small perturbations on inputs. A small perturbation on an input is termed an Adversarial Example (AE) if it leads to a different prediction from the label assigned to the original input. Over the past decade, numerous studies have attempted to frame DL robustness evaluation as a binary or worst-case problem, addressing questions like “if AEs exist (in a small local region of the original input)?” or “what is the closest AE to the original input?” [62, 6, 22]. Recently, emerging studies are adopting a probabilistic view, formulating robustness verification as a statistical inference problem [56, 51, 66, 50, 21, 8, 7]. Such a probabilistic robustness notion is arguably of more practical interest than binary/worst-case ones, because it provides an overall evaluation of how robust the model is whenever an AE can be found [56, 21] and accepts residual risks that are more realistic to achieve [51, 66]. We concur with this view and argue that T2I DMs also necessitate probabilistic robustness verification, which, to the best of our knowledge, is absent in SOTA.
@@
In this work, we first formally define the probabilistic robustness of T2I DMs against stochastic perturbation; then, we devise an efficient framework, named ProTIP, to evaluate and verify it (assuming a verification target is given). In addressing the statistical inference problems within ProTIP, we encounter new challenges specific to the generative nature of DMs. These challenges, distinct from traditional DL tasks like classification or regression, include: i) the generation process is very computationally insensitive; ii) the identification of a perturbed input as an AE entails comparing distributional differences, which is fundamentally harder compared to DL tasks like classification where an AE is identified by a change of the prediction label. To tackle the challenges, we employ sequential analysis with efficacy and futility early stopping rules in the statistical testing for indicating AEs, and adaptive concentration inequalities to dynamically determine the “just-right” number of stochastic perturbations whenever the verification target is met. Experimental results validate the effectiveness and efficiency of our ProTIP, while an application of ProTIP is also presented to showcase its use in ranking common defences against text perturbations.
In summary, key contributions of this paper include:
@@
  • Problem formulation: For the first time, we formulate the probabilistic robustness verification problem for T2I DMs against stochastic perturbation.
@@
  • Efficient solution: To solve the formulated problem, we develop an efficient framework, ProTIP, which incorporates several sequential analysis methods to dynamically determine the sample size and thus enhance the efficiency.
@@ 2.1 Text-to-Image Diffusion Models
Generative AI has been thriving in the multi-modal field, with DMs emerging as a powerful new family of deep generative models with record-breaking performance in many applications, including video generation [59], image reconstruction [49], and T2I generation [3]. DMs are a class of probabilistic generative models that apply a noise injection process, followed by a reverse procedure for sample generation. Importantly, DMs allow a notion of control at generation time through so-called guidance, enabling several complicated generative tasks, among which T2I DMs are examples of guided image generation tasks using descriptive text as the guidance signal. Models like Stability AI’s Stable Diffusion [42] and Google’s Imagen [44], trained on large-scale datasets of annotated text-image pairs, can generate high-quality images based on simple natural language descriptions. Commercial products like DALL-E 3 [3] and Midjourney [1] have demonstrated remarkable capabilities in diverse text-to-image applications, enriching the field. We refer readers to Appendix A for more details on T2I DMs.
@@ 2.2 Deep Learning Robustness
DL models are notoriously unrobust to small perturbations [62, 6, 22]. While definitions of robustness vary in literature, they share a common intuition that a DL model’s decision should remain invariant against small perturbations on a given input—typically it is defined as all inputs in a region $\eta$ have the same prediction label, where $\eta$ is a small norm ball (in a $L_{p}$-norm distance) of radius $\gamma$ around an input $x$. A perturbed input (e.g., by adding noise on $x$) $x^{\prime}$ within $\eta$ is an AE if its prediction label differs from $x$.
In the Safe AI community, DL robustness has been the property in the spotlight. Many studies on evaluating DL robustness have been done, framing the problem in different ways. In Fig. 2, we summarise 4 common ways of formulating the problem, inspired by [8]. Earlier works, such as [13, 24, 43], formulate the verification problem as a binary question by asking if any AEs can be found within a given input norm-ball of a specified radius, cf. Fig. 2(a). Such “binary robustness” can be normally evaluated in two ways [24, 43, 60]: either through reachability algorithms that aim to determine the lower and upper bounds of the output within the input norm-ball $\eta$, using layer-by-layer analysis; or by solving it using SAT/SMT solvers as a variety of constraint-based programming problems. Fig. 2(b) poses a similar yet different question: what is the maximum radius of $\eta$ such that no AE exists within it? Intuitively, it is finding the “largest safe perturbation distance” for input $x$ [2, 35, 57, 58]. While in Fig. 2(c), it evaluates the model’s robustness by introducing adversarial attacks to cause the maximum prediction loss in the specified norm ball $\eta$. It is often applied in adversarial training to enhance the model’s robustness to resist attacks [33, 53].
@@
The three aforementioned methods all estimate the robustness of the model by detecting the presence of AE or the AE that gives the maximum loss/safe-radius, the so called deterministic robustness [66]. As argued by [56], they suffer from two major drawbacks: i) they fail to convey how robust the model is whenever an AE is found; ii) they pose scalability challenges when the model is large. Thus, recent works [56, 51, 66, 50, 21, 8, 7] develop a new probabilistic view, by defining robustness as the proportion of AEs inside the norm-ball $\eta$, cf. Fig. 2(d). This probabilistic notion is arguably more practical, because: i) binary/worst-case robustness focusing on extreme cases is neither necessary nor realistic, especially when the model is large; knowing the proportion of AEs is more relevant; ii) since all practical applications have acceptable levels of risk, it suffices to demonstrate that the violation probability is below a required threshold, rather than confirming it to be exactly zero. Without loss of generality (WLOG), we illustrate probabilistic robustness using a DL classification task as:
@@ 2.3 Robustness of Text-to-Image Diffusion Models
Despite variations of definitions, the key to evaluating and improving the model’s robustness is detecting AEs. The approaches to this end are often referred to as “adversarial attacks” via input perturbations. Although such perturbations are commonly referred to as “attacks”, they are not necessarily malicious actions of attackers [68]. They may also represent natural sensor white noise [17] or benign human errors that follow a stochastic generation process. Note that in this work, we adopt such a generalised terminology of AEs to represent small and random perturbations. For T2I models, such perturbations can be classified into character-level, word-level, sentence-level or multi-level, depending on the granularity of input perturbations. Recent studies [34, 10] show that T2I DMs are very sensitive to black-box attacks like text perturbation. In [11, 69], T2I DMs are shown to be vulnerable to realistic human errors (e.g., typos, glyphs, phonetic errors), exposing significant robustness issues due to weak text encoders.
These studies focused only on deterministic robustness (e.g., maximised prediction loss). To the best of our knowledge, there is no dedicated exploration into the probabilistic robustness of T2I models when they are subject to stochastic perturbations, and our ProTIP is the first verification framework for this problem, backed by statistical guarantees. Moreover, ProTIP addresses unique challenges arising from the generative characteristics of T2I DMs, by adopting sequential analysis and adaptive concentration inequality to improve the verification efficiency, details of which are provided in the next section.
@@ 3.1 Problem Statement
where $T$ denotes the CLIP [41] model’s text encoder, $s$ is a similarity measurement function (e.g., cosine similarity), $\gamma$ denotes a given threshold on similarity. While $I$ is an indication function as defined in Def. 1, its value now depends on whether the output distributions before and after the perturbation differ. $Pr(x^{\prime})$ indicates the probability that $x^{\prime}$ is the next perturbed text generated randomly.
@@ 3.2 Stochastic Perturbation Generation
While ProTIP can cope with any text perturbations generated stochastically, WLOG, we study five character-level text perturbation methods from [36]. First, we set the perturbation rate like 10%, indicating that 10% of the words in the sentence will be perturbed. Then, one of the five perturbation methods (insert, substitute, swap, delete, and keyboard error) shown in Table 2 of Appendix 0.B will be randomly selected. Using insertion perturbation as an example, a letter from the 26 English alphabets is randomly selected and inserted into a randomly chosen position within a word in the sentence. This process is repeated until the perturbed sentence reaches a 10% perturbation rate, ensuring it’s not a duplicate before stopping. Taking “A white dog plays with a red ball on the green grass” as an example, after insertion perturbation, we get the perturbed sentence “A white daog plays with a red ball on the green grass”. More details and examples about each text perturbation method can be found in Table 2 of Appendix B
@@
As Def. 1, the intuition behind robustness suggests that the outputs of a DL model should remain invariant to perturbations on the inputs, while the inputs, before and after the perturbation, should be deemed very similar—thus, Def. 1 restricts the perturbation in a norm-ball $B(x,\gamma)$ with a small radius $\gamma$. Such perturbation distance control is easily achievable in Computer Vision where the inputs (either raw pixel values or the latent space feature values) can be normalised and projected into a continuous input space. However, for DL models take text inputs (like T2I DMs), the discrete nature of text make it more challenging to study robustness [29] as a single character change may completely alter the semantics meaning. E.g., “a red ball on …” and “a red bell on …” have totally different semantic meanings with only one character perturbed. Thus, to study the robustness of T2I DMs, first we need to ensure that the perturbed text $x^{\prime}$ has the same semantics to the original text $x$. Following the idea from [11, 69, 64, 31, 47, 63] for the same problem, we employ CLIP [41] as the model for text encoding, leveraging CLIP’s ability to represent image and text information while preserving their relationships.
Therefore, as per Def. 2, we use CLIP’s text encoder $T$ to extract the original embeddings $T(x)$ and perturbed sentence embeddings $T(x^{\prime})$ to compute the similarity score $s(T(x),T(x^{\prime}))$ (e.g., by cosine similarity). Only when the similarity score is greater than a given threshold $\gamma$, we deem $x^{\prime}$ is a valid perturbation, where $\gamma$ can be determined by the method developed in [29] (which yields the maximum perturbation distance in the continuous embedding space that preserves the semantics meaning for a given text). In this way, we ensure that only valid perturbations (sharing similar semantics as the original text) are used for robustness evaluation in ProTIP.
@@ 3.3 Indication of Adversarial Examples
which corresponds to the cosine similarity between visual CLIP embedding $V(y)$ for an image $y$ and textual CLIP embedding $T(x)$ for a caption $x$. We refer readers to Appendix B for details on the CLIP Score. For each image $y^{\prime}_{i}$, we calculate a set of i.i.d. “CLIP scores” $\{c^{\prime}_{1},\dots,c^{\prime}_{n}\}$ by measuring $c^{\prime}_{i}:=CLIPScore(x,y_{i}^{\prime})$. A CLIP score $c^{\prime}_{i}$ represents the coexistence likelihood of the image $y^{\prime}_{i}$ and the original text $x$. We repeat the same calculation for each image $y_{i}$ generated from the original text $x$, collectively forming another set of CLIP scores $\{c_{1},\dots,c_{n}\}$.
@@
  • $\mathcal{H}_{1}$: The CLIP scores of the $c^{\prime}_{i}$ group is smaller than the $c_{i}$ group.
While there are established statistical testing methods, e.g., the t-test, u-test, and their variants [27], that can be employed to answer the above question, their applicability depends on whether the data distribution meets certain assumptions. Our ProTIP is compatible with any applicable statistical tests in this step, cf. later Sec. 4 for our choice in the experiments. We can now implement the AE indicator function in Def. 2 as:
@@
In contrast to conventional DL tasks like classification, where detecting an AE involves an “one-off” test to see if the prediction label changes, determining if a perturbation affects the T2I DM requires multiple generations and comparing the distributional differences between two sets of images generated before and after the perturbation. This shift in the evaluation approach leads to a significantly increased computational workload. That said, we employ sequential analysis with early stopping rules [23, 39, 54] for the hypothetical test in Eq. (5). Instead of collecting all data with the maximum sample size (fixed at the design time) and then analysing the data a single time by the end, sequential analysis conducts interim analyses during data collection, so that we can prematurely stop data collection at an interim analysis upon rejecting or accepting the null hypothesis. Thanks to early stopping, sequential designs will, on average, require fewer samples [27, Chap. 10].
In ProTIP, we adopt sequential analysis with both Efficacy and Futility stopping rules, by controlling the $\alpha$ (Type I error, false positive when a null hypothesis is incorrectly rejected) and the $\beta$ (Type II error, false negative when the null hypothesis is accepted and it is actually false), respectively. Intuitively, the Efficacy stopping rule says, if the analysis reveals a statistically significant result at some interim stage, data collection can be terminated because we reject $\mathcal{H}_{0}$ when observing the p-value $p<\alpha_{k}$ (where $\alpha_{k}$ is the cumulative Type I error rate at the $k$-th interim analysis, controlled by alpha-spending functions [27]). On the other hand, we may also stop for futility, which means it is either impossible or very unlikely for the final analysis to yield $p<\alpha$. This can be implemented by controlling the Type II error rate across interim analysis using a beta-spending function. We refer readers to [27, Chap. 10] for more details on the topic.
Specifically, we employ the R package rpact [55] to perform the sequential analysis by setting Pocock type alpha/beta spending functions [15] to determine the sample size and the thresholds for efficacy/futility early stopping rules. Details are presented in Sec. 4 and Appendix D